OpenStudy (anonymous):

Derivative of a natural log function?? Answer provided Find the derivative: f(x) = ln(sqrt(4+x^2)/x) The answer is supposed to be -4/(x(x^2+4))

hartnn (hartnn):

\(\sqrt A = A^{1/2}\) use this, so that you can use the property of log, \(\log A^B = B \log A\) which would simplify your expression.

OpenStudy (anonymous):

Given \(\ln u(x)\), you have \(\dfrac{d}{dx}[\ln u(x)]=\dfrac{\frac{d}{dx}[u(x)]}{u(x)}\)